OpenStudy (anonymous):

Solve for x: -3|x + 7| = -12 x = 5 over 3, x = −19 over 3 x = −3, x = −11 x = −3, x = 11 No solution

OpenStudy (anonymous):

please help ihave to catch up before its too late

OpenStudy (shamil98):

-3|x+7| = -12 First divide both sides by -3 |x+7| = 4 , |x+7} = -4 x + 7 =4 x + 7 = -4 x = -3 , x =-11

OpenStudy (jdoe0001):

\(\bf -3|x + 7| = -12 \implies |x + 7| = \cfrac{-12}{-3}\\ \quad \\ |x + 7| = \cfrac{12}{3}\implies \begin{cases} +(x + 7) = \cfrac{12}{3}\\ \quad \\ \bf -(x + 7) = \cfrac{12}{3} \end{cases}\)

OpenStudy (anonymous):

thank you a lot,guys i never payed attention in school and now im payin for it but im like far behind being far behind right now smh:(

OpenStudy (anonymous):

The minimum and maximum temperature for a day in Cupcake Town can be modeled by the equation below: 3|x − 8| + 15 = 18 What are the minimum and maximum temperatures for this day? x = 1, x = 17 x = −7, x = 9 x = 7 x = 9 No solution wut bout this